### v1

# Extending Quantum-Safe Communications to Real-World Networks: An Adaptive Security Framework

Abstract
The advent of quantum computing threats classical cryptographic mechanisms, demanding new strategies for securing communication networks. Since real-world networks cannot be fully Quantum Key Distribution (QKD)-enabled due to infrastructure constraints, practical security solutions must support hybrid operation. This paper presents an adaptive security framework that enables quantum-safe communications across real-world heterogeneous networks by combining QKD and Post-Quantum Cryptography (PQC). Building upon a hierarchical key management architecture with Virtual Key Management Systems (vKMS) and a centralized Quantum Security Controller (QuSeC), the framework dynamically assigns security levels based on node capabilities. By transitioning between pure QKD, hybrid, and PQC modes, it ensures end-to-end quantum-safe protection regardless of the underlying node capabilities. The framework has been implemented and validated on a Kubernetes-based containerized testbed, demonstrating robust operation and performance across all scenarios. Results highlight its potential to support the gradual integration of quantum-safe technologies into existing infrastructures, paving the way toward fully quantum-safe communication networks.
I INTRODUCTION
The development of quantum computing introduces a paradigm shift in the field of cybersecurity. While quantum computers promise revolutionary advancements in several fields, they also pose a significant threat to classical cryptographic algorithms. In particular, widely used public-key algorithms such as RSA and ECC are considered vulnerable to quantum attacks, potentially compromising the asymmetric primitives that underpin key exchange and digital signatures in modern communication systems [easttom2022quantum]. This renders current mechanisms for establishing and managing cryptographic keys insecure.
To address these threats, two primary approaches have been developed for achieving quantum-safe key establishment. The first approach is Post-Quantum Cryptography (PQC), which replaces classical public-key algorithms with alternatives based on hard mathematical problems believed to resist quantum attacks. PQC algorithms can be implemented using existing classical hardware infrastructure, making them easily deployable across current networks. However, their security ultimately depends on computational assumptions [jenefa2023pqc].
The second approach is Quantum Key Distribution (QKD), which avoids public-key cryptography entirely by generating symmetric keys between two legitimate parties using the principles of quantum mechanics. Unlike PQC, QKD provides Information-Theoretic Security (ITS), guaranteed by physical laws rather than computational hardness assumptions [nadlinger2022experimental].
However, several challenges hinder its widespread adoption within current communication infrastructures. First, QKD requires specialized quantum hardware, which increases deployment cost and complexity [diamanti2016practical]. Second, the achievable distance is limited by the lack of quantum repeaters, restricting QKD to metropolitan-scale links. Third, extending QKD beyond point-to-point connections often requires trusted relay nodes, which may weaken the security guarantees by introducing additional points of vulnerability. Fourth, many QKD systems require dedicated dark fibers for quantum signal transmission, complicating integration with existing optical infrastructures and limiting scalability.
Although a fully QKD-enabled network represents the ideal quantum-safe communication paradigm, current constraints make such deployments impractical today. Instead, real-world networks are usually heterogeneous, comprising both QKD-enabled and classical nodes that rely on PQC, as depicted in Fig. 1. Each node is defined as a site that may include one or more pieces of network equipment, quantum equipment, or other related infrastructure. Hybrid architectures that strategically integrate both approaches offer a realistic path forward: they enable the gradual incorporation of QKD while ensuring continuity of service through PQC. This heterogeneity, however, introduces new challenges for providing end-to-end (E2E) quantum-safe communication across mixed networks.
Figure 1: A representation of a heterogeneous network comprised of QKD-enabled and classical nodes.
Building on our previous work on scalable QKD architectures [sanz2025quantumsafescalablenetworksopen], which introduced a hierarchical Key Management System (KMS) with Virtual KMSs (vKMS) and a centralized Quantum Security Controller (QuSeC) for efficient KMS discovery and multi-hop relay path computation, this paper extends the model to heterogeneous networks. While that earlier solution improved scalability with compliance with ETSI specifications, it assumed universal QKD capability, limiting its applicability in transitional or hybrid environments.
To address this gap, we propose an adaptive security framework that unifies QKD and PQC under centralized management. The framework introduces a layered architecture with vKMSs and a QuSeC that dynamically assigns one of four security levels, ranging from pure QKD to pure PQC, with hybrid strategies that leverage Key Derivation Functions (KDFs). This ensures that each communication session receives the strongest feasible quantum-safe protection, regardless of node capabilities. Therefore, the main contributions of this work are the following:
  1. 1.
    Adaptive Security Levels: Definition of four security levels that integrate QKD and PQC flexibly, enabling quantum-safe protection in heterogeneous networks.
  2. 2.
    Unified Architecture: An adaptive security framework that extends quantum security to heterogeneous real-world networks by combining QKD and PQC under centralized management, ensuring tailored security for all nodes. The framework also applies communication network concepts to operate in a efficient, scalable and resilient form. The framework leverages communication network principles to operate efficiently, scalably, and resiliently. It also extends the vKMS + QuSeC architecture to orchestrate security policies, assignments, and interoperability across QKD and PQC domains.
  3. 3.
    Standard-aware: A standard-aligned model compatible with ETSI specifications, supporting incremental integration of QKD technologies into existing infrastructures without disrupting operation.
The rest of the paper is organized as follows: Section II reviews prior work on QKD and PQC integration, and on extending quantum-safe security to heterogeneous networks. Then, Section III summarizes our previous work on E2E quantum key establishment in QKD networks, on which this proposal builds. Subsequently, Section IV introduces the adaptive security framework, describing the system architecture and its operation. Section V presents the implementation, testbed setup, and validation results. Finally, Section VI presents the main conclusions of the work.
II RELATED WORK
The combination of QKD and PQC for generating hybrid secure keys has attracted significant attention in recent years. The motivation of the hybrid approach is twofold: extending the range and applicability of QKD by complementing it with PQC mechanisms, and leveraging hybridization to obtain robust security guarantees, where the final session key remains secure as long as at least one of the contributing components is uncompromised.
Some theoretical foundations of hybrid Key Encapsulation Mechanisms (KEMs) are formalized in [bindel2019hybrid], where the security of several combiners, such as XOR-then-MAC (xTM) [giacon2018kem], dual-PRF [bellare2024symmetric], and the N Nested Dual-PRF, is introduced and proved. Their results can be used as the foundation for later works that integrate multiple key sources into secure session keys.
Building on this work, [ricci2024hybrid] proposes a three-key combiner that takes as input 256-bit keys from different sources: a classical key generated via Diffie-Hellman, a PQC key generated with KYBER, and a QKD key. They implement and validate the proposal in an FPGA platform, and they also prove IND-CCA security. Similarly, [aquina2024quantum] introduces a quantum secure communication system leveraging a PRF-then-XOR split-key combiner, showing that security can be guaranteed even if only one of the underlying primitives remains secure.
Several works have demonstrated the feasibility of such combiners in realistic cryptosystems. For instance, [zeng2024practical] defines a protocol for symmetric key establishment in QKD–PQC networks between two nodes, considering that several key distribution channels are available between them. The final key is derived through a hybrid KDF applied to multiple distribution channels. The obtained results validate the hybrid protocol where QKD and PQC interoperate within a joint quantum-classical network. In addition, [garms2024experimental] presents Muckle++, a modular cryptosystem integrating hybrid authentication and hybrid key exchange. The proposed protocol derives the secure keys from combining QKD, PQC, and currently used public-key cryptography using a HMAC-based KDF (HKDF). The protocol is validated through real QKD hardware, FPGA-based PQC-KEM acceleration on an FPGA, and server-side implementations. Along similar lines, [doring2022post] integrates PQC key exchange mechanisms into a QKD platform, enabling interoperability between the two technologies and providing a means to deliver quantum-safe keys to nodes where QKD is not available. They demonstrate the feasibility of the proposed scheme in the Berlin OpenQKD testbed, enabling quantum-safe communication between network nodes and extending the solution to wireless, last-mile, and end-user applications. These contributions highlight that hybrid protocols are both practical and deployable, bridging the gap between theoretical combiners and operational systems.
The role of KMSs in hybrid architectures has also been explored in detail. For instance, [james2023key] provides a comprehensive analysis of KMS requirements for large-scale QKD networks, classical and quantum KMSs and emphasizing the need for E2E guarantees even in the presence of compromised trusted nodes. In addition, to strengthen the overall security guarantees provided by a QKD network, they leverage PQC-based relaying of QKD keys between end nodes, with the final session key derived via hybridization at the KMS level. In a complementary direction, [brauer2024linking] demonstrates several approaches for achieving quantum-safe key exchange between three QKD testbeds in Europe, Berlin, Madrid, and Poznan, leveraging PQC-based encryption to emulate long-range QKD links. Different key exchange methods were defined to enable E2E communication between all the nodes of all the participating networks. In addition, they define a key hybridization process by applying a KDF as an internal process on the KMS that manages different internal key stores. These works illustrate how KMSs can serve as the central point for orchestrating hybridization in large-scale deployments.
These efforts primarily focus on extending or strengthening fully QKD-enabled networks and providing a solid foundation for the combination of several input keys through KDFs. However, some works also address environments where only a subset of nodes are QKD-capable. In this context, [zhu2023qkd] introduces the concept of partially-QKD-deployed optical networks (pQKD-ON), where only part of the optical nodes are deployed with QKD systems, and define the problem of key provisioning for secure E2E services in such networks. This work highlights the core challenge of enabling secure E2E communication in heterogeneous settings. They propose a multi-level key pool slicing scheme where QKD key pools are sliced and some portions of keys are pushed outward via lightweight secure transfer. This approach enables having pre-shared QKD key-pairs between non-QKD-deployed nodes. Although they address key provisioning in heterogeneous networks, their approach lacks the centralized management and dynamic security level adaptation that our framework provides. Their solution focuses on static key pool slicing without the unified control plane and or standards compliance that enables real-world deployment.
In summary, prior work has provided a strong foundation on secure key combiners for hybridization, practical hybrid protocols validated with QKD and PQC, and KMS architectures that support multi-domain integration. However, these efforts largely target homogeneous or fully QKD-enabled networks, not addressing the broader problem of dynamically extending quantum-safe security to heterogeneous environments where some nodes lack QKD capability. To the best of our knowledge, no existing framework provides a unified mechanism for adapting the security level depending on the node capabilities in heterogeneous and real-world networks. Our proposed framework addresses this gap by combining QKD and PQC under centralized management to deliver tailored, quantum-safe protection for all nodes, regardless of their underlying capabilities.
III BACKGROUND: E2E QUANTUM KEY ESTABLISHMENT IN QKD NETWORKS
This work builds on our previous work on scalable QKD architectures, which presented a foundational hierarchical E2E key establishment framework based on a centralized, programmable architecture leveraging Software-Defined Networking (SDN) principles. This work addressed the challenges of KMS identification, relay path discovery, and scalability in QKD networks. The proposed architecture introduced two key components that form the basis for the current work: the Virtual Key Management System (vKMS) and the Quantum Security Controller (QuSeC).
On the one hand, the vKMS operates as an abstraction layer within each QKD-enabled node, managing multiple underlying KMS instances associated with different QKD modules. The vKMS presents a unified interface to applications, hiding the complexity of the QKD infrastructure and enabling efficient key retrieval and relay coordination. This abstraction enables applications to request quantum keys without needing knowledge of the underlying topology or KMS configuration.
On the other hand, the QuSeC operates as a centralized management entity that maintains a global view of the QKD network topology. Leveraging SDN principles, the QuSeC computes optimal E2E paths for key distribution, considering different metrics such as available key rate, hop count, or physical distance. It coordinates the installation of relay rules across multiple KMS instances and manages the complete lifecycle of multi-hop key relay operations.
This hierarchical design supports both direct key delivery within direct QKD endpoints and multi-hop trusted relay across multiple nodes. The framework ensures compliance with ETSI specifications, including ETSI GS QKD 014 [ETSI_GS_QKD_014] for key delivery, ETSI GS QKD 015 [ETSI_GS_QKD_015] for SDN integration, and the ETSI GS QKD 020 [ETSI_GS_QKD_020], which is under development, for KMS interoperability in relay procedures.
While this work improved scalability and provided a standard-aware key delivery mechanism for QKD networks, it assumed universal QKD capability across all network nodes, limiting its applicability to purely QKD-enabled environments. The current work addresses this limitation by extending the vKMS and QuSeC architecture beyond QKD-only deployments, enabling adaptive security management across heterogeneous networks where some nodes lack QKD infrastructure.
IV PROPOSED FRAMEWORK
This section presents an adaptive framework that extends quantum-safe protection to all nodes within a network. The framework employs dynamic security assignment, adapting the level of security to the specific characteristics and capabilities of each node. In this context, the centralized network and security management plays a crucial role, providing a comprehensive view of the network to enable efficient and robust security management.
The framework seamlessly integrates QKD and PQC to enable quantum-safe E2E key delivery across heterogeneous network environments. In addition, it includes a secure and efficient key provisioning mechanism allowing nodes without direct access to QKD links to participate in quantum‑safe sessions via relay and hybridization strategies. As a result, communications maintain resistance to quantum adversaries even when only a subset of the network is QKD‑enabled.
IV-A System Architecture
Figure 2: System Architecture of the proposed framework.
The proposed architecture comprises two different node types. On the one hand, QKD-enabled Nodes (QN) are defined as nodes with access to QKD infrastructure capable of generating and exchanging QKD keys with peer QNs. Each QN hosts at least one KMS responsible for secure storage, synchronization, and delivery of QKD keys on demand. In Fig. 2, nodes D, E, and F are categorized as QN.
On the other hand, Classical Nodes (CN) lack QKD infrastructure, but are capable of employing PQC algorithms to achieve secure communication. In the example architecture shown in Fig. 2, nodes A, B, and C represent the CN category.
Considering this, the proposed architecture is structured into several communication planes, as depicted in Fig. 2:
  • Data Plane: This plane comprises the network devices, such as the switches or routers, used for the user traffic transmission.
  • Quantum Plane: This plane comprises the physical QKD modules, including transmitters, receivers, and quantum links. In this plane, QKD keys are generated point-to-point between peer nodes and pushed to their associated KMSs.
  • Quantum Key Management Plane: This plane comprises the KMSs responsible for managing QKD keys. Each QKD module in the Quantum Plane is paired with a KMS. The main functions of the KMS include secure storage and management of QKD keys, on demand delivery of QKD keys to authorized applications, and synchronization with the peer KMS.
  • Security Management Plane: A vKMS at every node, whether QN or CN, abstracts cryptographic operations from applications and implements the mechanisms instructed by the controller to derive a quantum-safe symmetric key between communication endpoints. The vKMS extends the functionalities of traditional KMSs by supporting key relay through quantum-safe channels, secure key generation via KDF, or PQC-based key establishment.
  • Control Plane: This plane comprises the QuSeC, which is responsible for managing both the network and security operations. In addition to standard SDN controller functions such as path computation or resource allocation, the proposed system introduces a Security Level Assignment Function. This function dynamically assigns and adjusts the security mechanism for the requesting communications based on the node types of the requesting nodes.
  • End-User Plane: This plane includes the end-user applications that obtain quantum‑safe keys from the local vKMS for consumption by upper‑layer protocols.
This decomposition separates concerns: cryptographic material is never exposed to the QuSeC, applications are isolated from vendor‑specific KMS details, and policies are enforced uniformly across heterogeneous nodes.
IV-B Operational description of the security framework
The adaptive security framework ensures quantum-safe E2E communication by dynamically assigning Security Levels based on the involved nodes. The framework is designed to always provide the highest Security Level achievable given the network capabilities and node types involved in the communication.
Figure 3: Operational flow of the proposed framework for all levels.
The complete E2E process operates through six sequential steps that transform an application’s key request into a secure, quantum-safe session key. Fig. 3 depicts the high-level operational flow of the proposed framework, illustrating how these steps interact across the different architectural planes. The process begins when an application initiates a secure session request and concludes with the delivery of quantum-safe keys to both communicating endpoints.
  1. 1.
    Secure Session Request. An initiating application requests a quantum-safe communication session with another application to its associated vKMS. This request is made sending a Get Key message defined in the ETSI GS QKD 014 API. The vKMS receives this request and prepares to coordinate with the QuSeC to deliver the secure key.
  2. 2.
    Security Level Request. The source vKMS cannot independently determine the optimal security approach since this depends on the capabilities of both endpoints and the quantum path availability. Therefore, upon receiving the request, the vKMS queries the QuSeC by sending a security_level_request (APP_SRC | APP_DST) message to know the Security Level that corresponds to the communication.
  3. 3.
    Security Level Assignment. The QuSeC executes the Security Level Assignment Function to determine the best security level. It resolves endpoint capabilities (QN or CN) and path feasibility, selects one of the four Security Levels (detailed in Section IV-B1) and mechanisms, and responds to the vKMS with the assigned level.
  4. 4.
    Security Level Configuration. With the Security Level determined, the vKMS requests detailed configuration instructions from the QuSeC. The QuSeC then computes a comprehensive configuration policy with instructions tailored to the selected Security Level. These instructions range from minimal relay path installation to PQC or KDF configuration. The QuSeC distributes these configuration policies to all involved vKMSs and KMSs, programming the distributed system to execute the chosen security mechanisms.
  5. 5.
    Secure Key Derivation. This is the execution phase where all configured components work together to generate the shared session key. The specific operations depend on the assigned Security Level: QKD-enabled nodes may retrieve quantum keys from their KMSs, relay operations may forward keys through intermediate trusted nodes using OTP encryption, PQC operations may establish shared secrets using post-quantum KEMs, and hybrid approaches may combine multiple key sources through cryptographic KDFs. Throughout this process, each vKMS performs its assigned role while maintaining synchronization with its peer to ensure both endpoints derive the identical session key.
  6. 6.
    Secure Key Delivery. Once the quantum-safe session key has been successfully derived, each vKMS delivers the key to its respective application over a secure, authenticated channel. The applications can then proceed with their secure communication using the established session key.
The key identifier exchange between applications occurs after step 6, when the initiating application transmits the session key identifier to the target application through their application-level protocol. When the target application subsequently requests its copy of the session key using a Get Key with ID message, the process is streamlined: the QuSeC maintains session state and recognizes this as a session key retrieval rather than a new session establishment, and instructs the target’s vKMS to deliver the previously derived key directly, bypassing steps 4 and 5. The mechanism by which the key identifier is conveyed is out of the scope of this work.
After both applications have received the secure keys, they can proceed with communication using the established session key. The framework is designed to be protocol-agnostic, allowing the quantum-safe key to be integrated into different security mechanisms, such as pre-shared keys, session keys, or encryption keys, depending on the requirements of the underlying communication protocol.
The following subsections describe in more detail the specific procedures of Security Level Assignment, Security Level Configuration, and Secure Key Derivation for each of the four defined Security Levels.
IV-B1 Security Level Assignment.
For the assignment of a Security Level, each vKMS, upon receiving a secure key request from an application, queries the QuSeC to request a Security Level for the communication between the initiating and the target applications. The proposed framework adapts the Security Level based on whether the involved applications belong to QN or CN, and on the quantum path availability. The QuSeC assigns a Security Level to each E2E communication and coordinates all the associated vKMSs and KMSs to execute the required operations to derive a symmetric E2E key with the highest security possible. In this context, after analyzing all possible scenarios, four different Security Levels have been defined:
  • Security Level 1 - Direct QN-to-QN. This Security Level applies when two applications, each belonging to QNs directly connected via a QKD link, request secure communication. The assigned Security Level is ITS, as no classical key exchange is involved, thereby fully preserving the quantum-safe properties of QKD. Fig. 4(a) shows the high-level operation of Level 1, between applications on nodes E and F.
  • Security Level 2 - Multi-hop QN-to-QN. This Security Level is assigned when two applications on QNs without a direct QKD link require secure communication. In such cases, the QKD with a trusted relay approach is employed, where keys are forwarded through intermediate trusted nodes leveraging One-Time Pad (OTP) encryption. Within the declared trust boundary for relays and with secure control and API channels, secrecy arguments are information-theoretic for the key material. Fig. 4(b) illustrates this Security Level for applications on nodes D and F.
  • Security Level 3 - CN-to-QN. This level applies when one application resides on a node without QKD infrastructure (CN). In this case, a relay and hybridization approach is employed to ensure secure key generation. Specifically, a hardened KDF derives a session key from at least one QKD-derived segment and one PQC KEM shared secret. The resulting key remains secure as long as at least one input key is secure. This multi-source approach produces a quantum-resistant symmetric key by combining independent security mechanisms. Fig. 4(c) depicts the high-level operation of Level 3 between applications on nodes C and D, including the relay of a QKD key via node E and the hybridization process.
  • Security Level 4 - CN-to-CN. This Security Level applies when both communicating applications reside on CNs, meaning no QKD infrastructure is available. In this case, a purely PQC-based approach is employed to ensure secure key exchange and derivation. Additionally, a KDF hybridization approach can be leveraged by combining multiple cryptographic sources to further enhance resilience. While this level does not provide ITS as in QKD-based approaches, it ensures robust protection against both classical and quantum threats. Fig. 4(d) shows this Security Level for applications on nodes A and B.
In practice, selection follows a simple preference ordering: Level 1>Level 2>Level 3>Level 4Level\ 1>Level\ 2>Level\ 3>Level\ 4, subject to quantum path feasibility and node capacity.
Figure 4: High-level operation of each Security Level.
IV-B2 Security Level Configuration.
After level assignment, the vKMS sends a level-based request to the QuSeC, which computes a level-specific configuration that encodes the session context, the selected cryptographic suites, and the minimal set of instructions required by each participant. For Levels 1 and 2, this step has been described in detail in our previous work, and it primarily involves KMS identification and, when applicable, relay path installation. For Levels 3 and 4, however, additional structure is required.
In the case of Level 3, the configuration is built around three roles that reflect the functions of the participating vKMSs. The receiver role is for the vKMS at the classical endpoint, it is the only endpoint that will receive, decrypt, and unveil the relayed QKD key. It will also establish a PQC KEM-based key with the passive vKMS, and it will derive and deliver the final session key to its application. The passive role is for the vKMS at the QKD-enabled endpoint, and it is responsible for retrieving the QKD key from its KMS, establishing a PQC KEM-based secret with the receiver, and deriving and providing the session key to its own application. Between them, the relay role is for the intermediate KMS/vKMS pair at the QN that has a QKD link with the passive node. It performs the relay of the QKD key via PQC KEM encapsulation and OTP encryption to the receiver.
Although each participant has a distinct role, all of them receive the same session configuration message. This message consists of common session information, such as a unique identifier, along with a policy block specific to each participant. The policy block determines the actions to be performed, the selected cryptographic suites, and the identities of the endpoints involved.
The configuration for Level 4 is comparatively simpler. Here, both endpoints directly establish a shared secret using the configured PQC KEM. The QuSeC provides each endpoint with a bilateral set of instructions that specify the KEM suites to be used, and, when multiple input keys are present, the KDF recipe for combining them into the final session key.
IV-B3 Secure Key Derivation.
Once the configuration is installed, Secure Key Derivation proceeds directly according to the installed policy instructions. As previously mentioned, the procedures corresponding to Levels 1 and 2 have been introduced in the previous work. In summary, Level 1 consists of retrieving a QKD key from the KMS. Level 2 also involves retrieving a QKD key from the KMS, but adds a key relay procedure between different KMS instances, following the relay path configured by the QuSeC in the preceding step.
Figure 5: Operational flow of Secure Key Derivation in Level 3.
For Level 3, the role-scoped steps that must be performed are described as follows, and they are graphically represented in Fig. 5:
  1. 1.
    The receiver and the passive vKMS establish a shared secret by running the configured PQC KEM. In this process, the passive generates a public-private key pair and shares the public key with the receiver vKMS. The receiver vKMS then performs the encapsulation function using the passive’s public key to generate a shared secret and corresponding ciphertext. The ciphertext is then transmitted back to the passive vKMS, which performs the decapsulation function using its private key to recover the same shared secret, thereby establishing a PQC key between both endpoints.
  2. 2.
    The passive vKMS retrieves a QKD key from the KMS through the ETSI GS QKD 014 key delivery API via a Get key message, and transmits the corresponding key ID to the relay vKMS.
  3. 3.
    Upon receiving the key identifier, the relay vKMS retrieves the identical QKD key using the Get key with ID message of the ETSI GS QKD 014 API.
  4. 4.
    The relay vKMS transports the key using post-quantum cryptography: it performs the encapsulation function of the configured PQC KEM with the receiver’s public key to generate an auxiliary shared secret with the receiver and a ciphertext. The relay then applies OTP encryption between the auxiliary key and the QKD key. It transmits both the ciphertext and the encrypted key to the receiver vKMS.
  5. 5.
    The receiver vKMS performs the decapsulation of the received ciphertext using its private key to recover the auxiliary shared secret. It then performs the OTP decryption to recover the original QKD key.
  6. 6.
    Both the receiver and the passive vKMS apply the configured KDF using as input the KEM secret and the QKD key in order to derive the session key.
Figure 6: Operational flow of Secure Key Derivation in Level 4.
On the other hand, the configuration of Level 4 entails the following operations:
  1. 1.
    Both vKMS endpoints execute the configured KEM to generate a shared secret.
  2. 2.
    Optionally, they can incorporate dual KEMs or additional entropy sources if enabled by policy. In this case, all generated shared secrets are combined using the configured KDF to derive the final session key.
Fig. 6 depicts the operational flow for the key derivation in level 4, for a configuration where the QuSeC configured a policy with dual KEM enforcement.
V VALIDATION
This section presents the implementation used to validate and evaluate the performance of the proposed adaptive security framework. First, the testbed and implementation details are introduced. Next, the functional validation of the proposed security framework is presented, demonstrating its correct operation. Finally, performance evaluation results are discussed.
V-A Implementation and Testbed Setup
All system components were fully implemented, including the vKMS, KMS, and QuSeC modules, following the architectural design presented in Section IV and ensuring full compliance with ETSI GS QKD 014 specification.
The proposed solution was deployed on an experimental infrastructure consisting of several servers forming a virtualized node managed by an OpenStack orchestrator, providing more than 200 vCPUs, 500 GB of RAM, and 20 TB of storage. On this infrastructure, an experimental testbed that replicates the heterogeneous network topology depicted in Fig. 2 was deployed using Ubuntu 24.04 and Kubernetes orchestration. The testbed comprises six nodes: three Classical Nodes (CN A, B, C) and three QKD-enabled Nodes (QN D, E, F).
The most relevant implementation details for each component are as follows.
  • Each vKMS instance supports the complete ETSI GS QKD 014 API to interface with end applications, extended with mechanisms for communication with the QuSeC to receive Security Levels and the related instructions. They support adaptive security mechanisms, including PQC integration and hybrid key derivation, under policy-driven Security Level execution. Post-quantum operations were realized with standardized algorithms, specifically CRYSTALS-Kyber (ML-KEM) [FIPS203] for KEM operations.
  • The KMS instances also implement the ETSI GS QKD 014 API for interaction with the vKMSs for QKD key delivery, as well as the ETSI GS QKD 020 API for KMS interoperability in multi-hop QKD scenarios. In addition, they support the operations proposed in our previous work for communication with the QuSeC and the vKMS.
  • The QuSeC implements the Security Level Assignment Function and associated configuration tasks, including topology aware path computation for multi-hop QKD scenarios, and automated policy generation for each Security Level.
  • The QKD modules of the quantum plane have been implemented using SimulaQron, a distributed quantum network simulator designed for software development. However, as the framework relies on standardized ETSI APIs, it maintains full compatibility with real QKD hardware from any vendor. The simulator serves purely as a proof-of-concept platform, enabling comprehensive validation without requiring physical infrastructure.
V-B Functional validation
Comprehensive functional testing was carried out to validate the correct assignment of Security Levels across all communication scenarios. Table I summarizes the test cases executed, covering the four Security Levels. All test cases achieved 100% accuracy in Security Level Assignment, with the QuSeC correctly analyzing node capabilities and quantum path availability to determine the optimal configuration.
TABLE I: Test cases
Test Source App Target App Expected Level Assigned Level Result
T1 APP_E APP_F Level 1 Level 1 Pass
T2 APP_D APP_F Level 2 Level 2 Pass
T3 APP_C APP_D Level 3 Level 3 Pass
T4 APP_A APP_B Level 4 Level 4 Pass
In addition, for each Security Level, the complete E2E communication flow was validated, from initial application request to final key delivery.
  • Level 1 Validation: Direct QKD key retrieval between adjacent QKD nodes (E-F) completed successfully with information-theoretic security preservation. The vKMS instances correctly interfaced with the appropriate KMSs, KMS_E2 and KMS_F, via the ETSI GS QKD 014 API.
  • Level 2 Validation: Multi-hop QKD communication (D-F through trusted-relay node E) demonstrated proper relay path establishment and OTP-based key forwarding. The KMSs properly implemented the ETSI GS QKD 020 API for KMS interoperability for the QKD key relay.
  • Level 3 Validation: Hybrid CN-to-QN communication (C-D) validated the correct execution of the relay and hybridization mechanism: the QuSeC configured the appropriate roles and policies (Section IV-B2), and the protocol followed the six-step procedure defined in Section IV-B3, achieving successful PQC KEM establishment, QKD key relay, and final hybrid key derivation via the configured KDF.
  • Level 4 Validation: Pure PQC communication (A-B) demonstrated correct PQC-only operation: the QuSeC provided the policy configuration, and both vKMS instances executed the PQC KEM successfully, ensuring quantum-resistant protection in the absence of QKD infrastructure.
Overall, the implementation demonstrated full interoperability across heterogeneous node types and seamless integration between QKD and PQC. All key exchanges maintained compliance with ETSI specifications, while extending functionality to support adaptive security mechanisms.
V-C Performance Assessment
To evaluate the performance of the proposed framework, latency measurements were collected for each Security Level. First, complete E2E key establishment across all Security Levels was analyzed, which accounts for all operations required for both the initiating and target applications. The delay associated with transmitting the key identifier from the initiating to the target application was excluded, as it is application-specific and out of the scope of this validation.
In addition to the overall E2E latency, a detailed analysis was performed for each individual phase, Security Level Assignment, Configuration, and Secure Key Derivation, in order to quantify their individual contributions to total latency. Each test case defined in Table I was executed 100 times to obtain statistically significant measurements. The generated boxplots have been calculated with 98 percentile and show the median (horizontal line), the mean (x), the box (top limit 75 percentile and bottom limit 25 percentile), and the “whiskers” (that contain the 98% of results).
Fig. 7 shows the total delay for E2E key establishment. Level 1 demonstrates the fastest establishment, averaging 73.35 ms through direct QKD key retrieval, followed by Level 4 with 92.3 ms via purely PQC-based operations. In contrast, the more complex Security Levels exhibited higher latencies, with 170.38 ms for Level 2 and 154.97 ms for Level 3. Despite the performance differences across levels, all key establishment operations remained below 200 ms, confirming the framework’s suitability.
Figure 7: Complete E2E key establishment time.
Moving towards the first subprocess, Security Level Assignment was highly consistent across all scenarios, with assignment times ranging from 8 to 9 ms for both initiating and target applications. This uniformity demonstrates that QuSeC performs the decision-making process efficiently, contributing negligible overhead to total E2E time.
The Security Level Configuration phase for the initiating application, shown in Fig. 8, varied more significantly between levels. Level 2 exhibited the highest latency at 58.35 ms, due to the complexity of relay path computation and rule installation across multiple intermediate KMS nodes. This reflects the coordination complexity demands for multi-hop QKD scenarios. The remaining Security Levels maintain low configuration times, with average values ranging from 6 to 11 ms. These results indicate that configuration complexity in the proposed framework is primarily driven by QKD relay infrastructure rather than PQC operations.
Figure 8: Security Level Configuration time for initiating applications.
The final phase, Secure Key Derivation, showed the largest variation across Security Levels, as depicted in Fig. 9 for initiating applications. Level 3 exhibited the highest processing time, averaging 98.09 ms, the complexity of the six-step hybrid mechanism that includes PQC KEM establishment, QKD key relay, OTP operations, and KDF processing. Level 2 followed with 53.30 ms, primarily due to multi-hop relay operations and OTP-based key forwarding. Conversely, Level 4 and Level 1 achieved significantly lower derivation times, 31.31 ms and 17.47 ms respectively, owing to their streamlined processes.
Figure 9: Secure Key Derivation time for initiating applications.
Throughout all phases, target application procedures consistently demonstrate lower delays than those of the initiating application. This is a direct result of the QuSeC’s session state management, which prevents redundant processing for subsequent endpoints. As represented in Fig. 10, the target application’s contribution to total E2E time remains stable across all Security Levels, at approximately 30-35 ms, while the initiating application’s contribution varies according to the operations required by each subprocess.
Figure 10: Contribution of initiating and target applications to the complete E2E key establishment time.
These results validate the framework’s principle of adaptive security assignment. Measured latencies demonstrate that enhanced security features, from Level 1’s information-theoretic guarantees to Level 3’s hybrid quantum-safe protection, can be achieved without excessive operational penalties. This performance evaluation therefore confirms the framework’s practical viability for deployment in heterogeneous quantum-safe networks.
VI CONCLUSIONS
This work addresses the need for quantum-safe communications in real-world heterogeneous networks by introducing an adaptive framework that integrates QKD and PQC under centralized management. Unlike existing approaches limited to fully QKD-capable environments, this proposal extends quantum protection to all nodes, ensuring robust protection regardless of infrastructure constraints. By defining four dynamic security levels, the system enables tailored protection, balancing information-theoretic guarantees with practical deployability. While the default behavior prioritizes maximum security, the framework can be extended to consider specific service requirements and can even deny service requests when minimum security requirements cannot be satisfied.
The implementation and validation results demonstrate the framework’s ability to deliver secure keys across all communication scenarios. All tested security levels achieved reliable E2E operation, while maintaining full compliance with existing ETSI specifications. Performance evaluation further confirmed the feasibility of the proposed approach, highlighting the potential of the framework for deployment in transitional infrastructures where QKD adoption is partial or incremental.
In summary, this work establishes a practical path toward hybrid quantum-safe networks. By unifying QKD and PQC into a standard-aware, adaptive model, the framework provides the flexibility needed to adopt evolving technologies.
ACKNOWLEDGMENT
This work was supported in part by the European Comission through the project GN5-2 HORIZON-INFRA-2024-GEANT-01-SGA in the Working Group 6 Task 1 - Technology, in part by the Spanish Ministry of Science and Innovation in the project EnablIng Native-AI Secure deterministic 6G networks for hyPer-connected envIRonmEnts (6G-INSPIRE) (PID2022-137329OB-C44), and in part by the Ministry of Digital Transition and Public Administration with European Union funds – Next Generation EU, under Component 16.R1 of the Recovery, Transformation, and Resilience Plan.